Paul,mpho and Clifford earned 2520 for painting a house.paul worked for 30 hours,phone worked for 50 hours and Clifford worked for 60 hours.the money should be divided in proportion to the number of hours each worked.by means of calculations , show how much did each get
1
Expert's answer
2020-06-07T16:00:25-0400

"\text{Paul worked for 30 hours , mpho worked for 50 hours and Clifford worked for 60 hours}\\[1 em]
\text{Total working hours} =30+50+60=140 \text{ hours}\\[1 em]
\because\text{Paul worked for 30 hours}\implies \text{the ratio of Paul }=\frac{30}{140}=\frac{3}{14}\\[1 em]
\therefore \text{Paul will get} \frac{3}{14} \times 2520 =\fcolorbox{red}{aqua}{540} \\[1 em]
\because\text{Mpho worked for 50 hours}\implies \text{the ratio of Paul }=\frac{50}{140}=\frac{5}{14}\\[1 em]
\therefore \text{Mpho will get} \frac{5}{14} \times 2520 =\fcolorbox{red}{aqua}{900} \\[1 em]

\because\text{Clifford worked for 60 hours}\implies \text{the ratio of Paul }=\frac{60}{140}=\frac{3}{7}\\[1 em]
\therefore \text{Clifford will get} \frac{3}{7} \times 2520 = \fcolorbox{red}{aqua}{1080} \\[1 em]
\text{Confirm the solution} 540+900+1080=2520"